Consider the following reaction: A(g)⇌2B(g) Find the equilibrium partial pressures of A and B for each of the following different values of Kp. Assume that the initial partial pressure of B in each case is 1.0 atm and that the initial partial pressure of A is 0.0 atm. Make any appropriate simplifying assumptions.

1.PA, PB, Kp= 1.4

2.PA, PB, Kp= 1.8×10^−4

3.PA, PB, Kp= 2.0×10^5
Answer #1

first to draw ICE table.

then write the equilibrium constant equation kp.

then using the given condition to find P​​​​​​A and P​​​​​​B

A Soln 2B initial ch x 1-2x Withing the equlibrium.cometent sexu Ko - trobe . ti-za) ( int given upeliu (-28) 1.4 0 + 4x 4x a2xiOS Date --I-I I Kp = 20x105 (1-200² - x It Y -40 = 2x10 x. yx?_ 2 x 10*x+d=0 ex10s + (2+105)² - 4X4XL 2x4 *=581006 Ø 50000